OpenStudy (anonymous):

cos2x-cos^2x=0 use the identities

OpenStudy (anonymous):

Do you know the double angle formula for cos2x?

OpenStudy (anonymous):

all integral pi values\[2\cos^2x-1-\cos^2x=0 \rightarrow cosx=+1 or -1\]

OpenStudy (anonymous):

cos^2x+sin^2x-cos^2x=0 Sin^2x=0

OpenStudy (anonymous):

DID YOU ASK ME (PHANI)

OpenStudy (anonymous):

doesnt my answer need to be zero?

OpenStudy (anonymous):

from -sin^2x=0 what do i do next?